OpenStudy (anonymous):

Find the center and radius of the circle whose equation is x^ 2 + y^ 2 - 6x - 2y + 4 = 0

Nnesha (nnesha):

\[\huge\rm x^ 2 + y^ 2 - 6x - 2y + 4 = 0\] write x's y's separately \[\huge\rm (x^2 -6x) +(y^2 -2y) +4 =0\] move 4 to the right side and apply "completing the* square method for left side "

OpenStudy (anonymous):

well my options are (3,1), r = (-3,-1), r = (3,1) r = 6 (9,1), r = 36 so im unsure as to which one it'd be @Nnesha

Nnesha (nnesha):

there are 3 4 step in order to reach t othe answer

Nnesha (nnesha):

s*

OpenStudy (anonymous):

would it be (9,2) + 36 ? @Nnesha

Nnesha (nnesha):

nope .... are you familiar with completing the square method ?

OpenStudy (anonymous):

no not at all @Nnesha

Nnesha (nnesha):

completing the square method divide middle term b by 2 for example in first parentheses \[\huge\rm (x^2 -6x)\] divide middle term by 2

Nnesha (nnesha):

\[\huge\rm (x + \frac{ b }{ 2 })^2\] after that take square of 3 and add that to the right side